OpenStudy (anonymous):

Convert 0.09% to a decimal

OpenStudy (radar):

Since it is already in decimal form, simply move decimal two places to the left (divide by 100)

OpenStudy (radar):

.0009%

OpenStudy (anonymous):

0.09% as a decimal is 0.0009

OpenStudy (anonymous):

0.09% = 9/100%=9/100/100=9/10000=0.0009

OpenStudy (radar):

I should of left the percent sign off. amyvincen12 is correct.